OpenStudy (anonymous):

In a circle, the circumference and diameter vary directly. Which of the following equations would allow you to find the diameter of a circle with a circumference of 154 if you know that in a second circle the diameter is 14 when the circumference is 44?

OpenStudy (jdoe0001):

"something" varies directly to "something else" "something" = "some value" * "something else" y = n * x what's the value "n" or "variation constant"? we dunno but we know that when the diameter is 14, the circumference is 44 d = diameter c= circumference thus c = n*d c = 44, d = 14 44 = n * 14 solve for "n" to find the variation constant value then plug it in the other circle's equation

OpenStudy (mathmale):

"In a circle, the circumference and diameter vary directly" is another way of stating the familiar formula \[\pi*d\] You could proceed from here merely by substituting the circumference C = 154 to determine the related diameter, d, or you could pretend you don't know that the proportionality constant is pi and solve C=k*d for k using the data pertaining to the second circle. Note that k = 44/14 comes out to approximately pi.